OpenStudy (anonymous):

what is the steps and answer to this 8=-(8+n)

OpenStudy (anonymous):

8 = -1 * (8 + n) 8 = -8 - n 16 = -n n = -16

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

can you tell me how you did that

OpenStudy (anonymous):

where did you get the -1 from